a proton is release 2.0x10^-11m from the center of a 6.4 x 10^-18 C fixed charge. What is the speed of the proton when it is 0.50 m from the charge (mass of proton: 1.674 x 10^-27 kg)
Answer #1

let

q = 1.6*10^-19 C (charge of proton)
Q = 6.4*10^-18 C
r1 = 2*10^-11 m
r2 = 0.5 m
m = 1.674*10^-27 kg

Apply conservation of energy

final total energy = initial total energy

k*Q*q/r1 = k*Q*q/r2 + (1/2)*m*v^2

9*10^9*6.4*10^-18*1.6*10^-19/(2*10^-11) = 9*10^9*6.4*10^-18*1.6*10^-19/0.5 + (1/2)*1.674*10^-27*v^2

==> v = 7.42*10^5 m/s